Saeed’s Bakery has a fire and Saeed loses some of his cost data. The bits of paper that he recovers after the fire provide the information in the following table (all the cost numbers are Rupees).

Q

AFC

AVC

ATC

 MC

100

1200

1000

2200

 

 

 

 

 

800

200

A

B

1500

 

 

 

 

 

900

300

400

900

1300

 

 

 

 

 

1300

400

300

C

D

 

 

 

 

 

E

500

240

1080

1320

 

Saeed asks you to come to his rescue and provide the missing data in the five spaces identified as ABCD, and E.


1
Expert's answer
2021-08-31T08:50:14-0400

a) A= average fixed cost (AFC).

Output= 20.

AFC = total fixed cost / output.

AFC can also be calculated by TFC ÷ Q.

AFC = TFC ÷ Q

TFC = AFC x Q

TFC = 120 x 10.

TFC = 1200

A = TFC / output

A = 1200/ 20

A = $60

b) B = Average Variable Cost (AVC)

Output = 20

AFC + AVC = ATC

AVC = AFC - ATC

AVC = $150 - $60

AVC = $90

B = $90

c) D = Average Total Cost (ATC)

Output = 40

ATC = TC / Quantity (Q)

Total Cost (TC) = ATC x Q

TC = 130 x 30

TC = 3900

Marginal Cost (MC) = change in TC / change in quantity

Change in TC = MC x change in quantity

Change in TC = 130 x 10, when Q=40 and Q = 30

TC = 3900 + 1300 = 5200, when Q = 40

ATC = 5200 / 40 = $130, when Q= 40

D = ATC = $130

D = $130

d) C = Average Variable Cost (AVC)

AVC = ATC - AFC

AVC = 130 - 30

AVC = $100

C = $100

e) E = Marginal Cost (MC)

Output = 40 to 50 units

MC = change in Total Cost (TC) / change in quantity(Q)

TC = ATC x Q


When output =40, TC = 130 x 40 = 5200

When output = 50, TC = 132 x 50 = 6600

Marginal Cost = (6600 - 5200) / 10 = $140

E = $140



A = $60

B = $90

C = $100

D = $130

E = $140